Hello... I need help this is a block question where you have to find the equation ok for the first step the # of blocks is 4 so the difference is 0, step 2 has 9 # of blocks and the difference is 5, step 3 has 16 blocks and a difference of 7, step 4 has 25 blocks and a difference 9, step 5 has 36 blocks and a difference of 11....... the 10th step has 121 blocks and a difference of 21 I think this is called the Nth step or increasing numbers please HELP
Answer by stanbon(75887)

1st term = 4
2nd term = 9
3rd term = 4^2=16
4th term = 5^2=25
....
10th term = 11^2=121
nth term = (n+1)^2
Cheers,
Stan H.